OpenStudy (anonymous):

Solve the following system: y = x + 3 3x + y = 19

OpenStudy (anonymous):

(7, 4) (−4, 7) (4, 7) (4, −7)

OpenStudy (anonymous):

@imqwerty

imqwerty (imqwerty):

frm 1st equation u knw taht y=x+3 put this value of y in 2nd eq u get 3x+(x+3)=19 3x+x+3=19 4x=16 x=4 we knw that y=x+3 we also knw that x=4 so put x=4 in y =x+3 y=4+3 y=7 so (4,7)

OpenStudy (anonymous):

thank you !! I have 2 more left can u help

imqwerty (imqwerty):

yes (:

OpenStudy (anonymous):

sorry i meant 3 left but i will make a new post and tag u

imqwerty (imqwerty):

lol ok :D